# Floquet-engineered moiré quasicrystal patterns of ultracold Bose gases in twisted bilayer optical lattices

### v4

@@ V Conclusions and Discussions
We numerically study the formation of novel moiré quasicrystal patterns of driven ultracold Bose gases in twisted bilayer hexagonal lattices. Dynamical evolutions of the IPR and the DW amplitude reveal the stage for the emergence of moiré quasicrystal patterns. The system undergoes a symmetry transition from lattice symmetry breaking to rotational symmetry emergence during the modulation, ultimately forming novel moiré quasicrystal patterns. Each moiré quasicrystal pattern generally persists for less than 1 ms before transforming into the next symmetrical configuration. This symmetry transition heralds a new quantum dynamical phase transition.
The spatial symmetry of patterns evolves temporally when the frequencies or amplitudes are modulated. Increasing the modulation frequency prevents the formation of moiré quasicrystal patterns while preserves the original lattice characteristics. In contrast, a larger modulation amplitude generates more complex symmetrical structures. Notably, a $D_{12}$ pattern emerges at $a_{m}=100a_{0}$ and $\omega=200$ Hz, closely resembling the moiré quasicrystal in twisted bilayer graphene. Furthermore, the momentum-space patterns also exhibit high rotational symmetries, consistent with the real-space patterns at specific evolution times.
@@
Acknowledgments. Our work is supported by the National Natural Science Foundation of China (Grant No. 11774093), the Natural Science Foundation of Shanghai (Grant No. 23ZR118700), the Program of Chongqing Natural Science Foundation (Grant No. CSTB2022NSCQMSX0585), and the Innovation Program of Shanghai Municipal Education Commission (Grant No. 202101070008E00099).

### v5

@@ V Conclusions and Discussions
We numerically study the formation of novel moiré quasicrystal patterns of driven ultracold Bose gasses in twisted bilayer hexagonal lattices. Dynamical evolutions of the MSD and total DW amplitude reveal the stage for the emergence of moiré quasicrystal patterns. In our weakly driven system, pattern formation is closely related to the momenta of collective modes. By quantifying the radial and angular DW amplitudes, we obtain the distributions of the emerging collective modes. We conclude that these new collective modes are coupled only radially and eventually cannot be decoupled. The system undergoes a symmetry transition from lattice symmetry breaking to rotational symmetry emergence during the modulation, ultimately forming novel moiré quasicrystal patterns. Each moiré quasicrystal pattern generally persists for less than 1 ms before transforming into the next symmetrical configuration. This symmetry transition heralds a new quantum dynamical phase transition.
The spatial symmetry of patterns evolves temporally when the frequencies or amplitudes are modulated. Increasing the modulation frequency prevents the formation of moiré quasicrystal patterns while preserves the original lattice characteristics. In contrast, a larger modulation amplitude generates more complex symmetrical structures. In particular, a $D_{12}$ pattern emerges at $a_{m}=100a_{0}$ and $\omega=200$ Hz, closely resembling the moiré quasicrystal in twisted bilayer graphene. Furthermore, the momentum-space patterns also exhibit high rotational symmetries, consistent with the real-space patterns at specific evolution times.
@@
Acknowledgments. Our work is supported by the Natural Science Foundation of Shanghai (Grant No. 23ZR1418700), the National Key Research and Development Program of China (Grant No. 2025YFF0515201), the Joint Funds of the National Natural Science Foundation of China (Grant No. U25D8014), the National Natural Science Foundation of China (Grant No. 11774093), the Program of Chongqing Natural Science Foundation (Grant No. CSTB2022NSCQ-MSX0585), and the Innovation Program of Shanghai Municipal Education Commission (Grant No. 202101070008E00099).